# Sampling-based Distributed Training with Message Passing Neural Network

### v1

@@ 1 Introduction and Related Work
While edge-based methods are accurate in modeling physical systems, their memory requirement scales with the number of edges in a mesh. Hence these methods do not scale for modeling realistic physical problems where the number of nodes can be substantially higher ($O(10^{5}-10^{6})$).  Bonnet et al. (2022b) attempted to solve this issue by randomly sampling a set of nodes and edges based on Nyström approximation (Li et al., 2020a) and memory constraints of the GPUs available, thus creating a sparse domain. The sampling method yields similar errors as modeling complete graphs. Despite many attempts, edge-based methods, especially MPNNs, remain memory-intensive, constrained by the current limits of single GPU memory. Strönisch et al. (2023) proposed an approach implementing multi-GPU utilization for MeshGraphNets (Pfaff et al., 2020). In their methodology, the domain is partitioned across various GPUs, enabling inter-node communication. However, this communication is limited to node-based features within the latent space, excluding interaction among edge-based features. The errors incurred during training and testing in the multi-GPU setup were notably higher compared to single GPU training (discussed in more detail in subsequent sections), highlighting the imperative need to devise methodologies for training edge-based graph models on multi-GPU setups without compromising accuracy.
We present a sampling-based distributed MPNN (DS-MPNN) that involves partitioning the computational domain (or graph) across multiple GPUs, facilitating the scalability of edge-based MPNN to a large number of nodes. This scalability holds significant practical value for the scientific community. We combine the idea of domain decomposition and message-passing among GPUs to enable the training of message-passing neural networks for physical systems without sacrificing accuracy.
@@
    We devise a method of training and inference on multiple GPUs for MPNN without sacrificing accuracy.
@@ 3 Methodology
We now turn to DS-MPNN, our framework devised to train edge-based graphs on many GPU systems. DS-MPNN incorporates communication strategies tailored for MPNN-based graph methodologies across multiple GPUs. In this framework, prior to the formulation of a graph kernel $\mathcal{G}=(V_{s},\mathcal{E})$, the computational domain $\Omega$ is partitioned into distinct subdomains and each subdomain is allocated to a GPU. Each GPU is allocated a distinct subdomain $\Omega_{r}\subset\Omega$. This arrangement entails dividing $\Omega$ into $n_{\text{procs}}$ (total number of GPUs available for parallelization) subdomains $\Omega_{r}$, with each subdomain featuring an extended overlap of radius $r$. This extended overlap is pivotal for ensuring comprehensive kernel construction at the interior edges of the domain, thereby circumventing the issues of incomplete kernel formation that can result in discontinuities in the predicted solutions. This approach is depicted in figure 1
@@ 4.1 Darcy Flow (Structured Data)
We explore an example utilizing a structured grid, specifically focusing on Darcy flow. This example parallels the approach used in the Graph Neural Operator (GNO) as discussed by  (Li et al., 2020a), adapted for a single GPU setup with a notable distinction: we update the edge attributes following each radius hop or message-passing step, and both edges and nodes are subjected to sampling. The experiment investigates the mapping $a\rightarrow u$, where $a$ and $u$ correspond to simulation parameters and output variables, respectively.
@@
Figure 3: Comparison of model performance between single GPU (S-MPNN), four-GPU (DS-MPNN4) setup, and GCN for a test sample of Darcy flow dataset.
Table 1 compares the performance of GCN and MPNN models on the Darcy flow dataset, using both single and 4-GPU setups. GCN under-performs relative to MPNN, highlighting the superiority of edge-based methods in PDE surrogate modeling, even for simpler datasets. This discrepancy is largely due to GCN’s omission of edge representation, a key feature in node-based models. Furthermore, the DS-MPNN model shows comparable results on both single and 4-GPU configurations. The minor variance in test error between these two setups might result from the segmentation of the computational graph across different domains. Figure 3 corroborates this observation, illustrating similar predictions from both DS-MPNN and S-MPNN configurations, while GCN significantly lags behind.
@@ 4.2.2 High-fidelity AirfRANS dataset
For the high-fidelity mesh analysis, we employ the standardized AirfRANS dataset as detailed by Bonnet et al. (Bonnet et al., 2022a). Our focus in this study is to show the effectiveness of DS-MPNN across various network hyperparameters rather than to achieve the best accuracy. Hence, we use a specific subset of the original AirfRANS dataset, termed as the ‘scarce’ dataset by the authors, that consists of 180 training samples and 20 test samples. The dataset encompasses a Reynolds number variation between 2 million and 6 million, and the angle of attack ranges from $-5^{\circ}$ to $15^{\circ}$. The input node attributes bear resemblance to the low-fidelity dataset 4.2.1 but with an expanded dimension of unit surface outward-pointing normal for the node $v_{i}^{l={0}}\in\mathbb{R}^{7}$, and the edge attributes are characterized by $e_{ij}^{(l)}\in\mathbb{R}^{10}$. The model output is denoted as $y_{i}\in\mathbb{R}^{4}$. Table 4 presents the test error, quantified as the Root Mean Square Error (RMSE) loss, where the dataset is sampled once. These results indicate that the DS-MPNN framework maintains its efficacy even with a substantial increase in the node count in the dataset, consistently performing well across different GPU configurations. In contrast, the GCN exhibits poorer performance under similar conditions. Additionally, an experiment involving a network configuration without inter-GPU communication, but with distinct domains distributed across GPUs, resulted in divergent training outcomes which are different from (Strönisch et al., 2023) where no communication model worked better. This underscores the critical role of communication between GPUs for maintaining model stability and performance. Table 4 shows that using the DS-MPNN framework with more GPUs increases training and inference speeds. This is because each GPU deals with fewer edges and nodes than it would if the entire domain was on a single GPU, which speeds up model execution. However, the speed increase is not linear. This is due to overlapping regions in each domain, along with the interior divided nodes, which increase the total number of nodes processed through the network, controlling the overall speed improvement.
@@ 5 Conclusions
We address the issue of scaling edge-based graph methods to larger numbers of nodes by introducing distributed training for message-passing neural networks (MPNNs). When combined with node sampling techniques, this distributed approach allows us to scale a larger number of nodes without losing accuracy as compared to the single-GPU implementation. We also show comparisons to the graph convolution networks (GCNs) and establish that edge-based methods like MPNNs outperform node-based methods like GCNs - This work opens up new avenues for the use of edge-based graph neural networks in problems of practical interest where the number of nodes can be impractically large, $\sim O(10^{5}-10^{6})$, for single-GPU memory limits.

### v2

@@ 1 Introduction and Related Work
While edge-based methods are accurate in modeling physical systems, their memory requirement scales with the number of edges in a mesh. Hence these methods do not scale for modeling realistic physical problems where the number of nodes can be substantially higher ($O(10^{5}-10^{6})$).  Bonnet et al. (2022b) attempted to solve this issue by randomly sampling a set of nodes and edges based on Nyström approximation (Li et al., 2020a) and memory constraints of the GPUs available, thus creating a sparse domain. The sampling method yields similar errors as modeling complete graphs. Despite many attempts, edge-based methods, especially MPNNs, remain memory-intensive, constrained by the current limits of single GPU memory. Strönisch et al. (2023) proposed an approach implementing multi-GPU utilization for MeshGraphNets (Pfaff et al., 2020). In their methodology, the domain is partitioned across various GPUs, enabling inter-node communication. However, this communication is limited to node-based features within the latent space, excluding interaction among edge-based features. The errors observed in the multi-GPU setup proposed by (Strönisch et al., 2023) were notably higher compared to their single GPU training. This highlights the imperative need to devise methodologies for training edge-based graph models on multi-GPU setups without compromising accuracy.
We present a sampling-based distributed MPNN (DS-MPNN) that involves partitioning the computational domain (or graph) across multiple GPUs, facilitating the scalability of edge-based MPNN to a large number of nodes. Here, ‘distributed’ implies partitioned spatial domains that are put on different GPUs which can be on same or different machines. This scalability holds significant practical value for the scientific community. We combine the idea of domain decomposition and message-passing among GPUs to enable the training of message-passing neural networks for physical systems with no or minimal loss in accuracy.
@@
    We devise a method of training and inference on multiple GPUs for MPNN with no or minimal loss in accuracy.
@@ 3 Methodology
We now turn to DS-MPNN, our framework devised to train edge-based graphs on many GPU systems. DS-MPNN incorporates communication strategies tailored for MPNN-based graph methodologies across multiple GPUs. In this framework, prior to the formulation of a graph kernel $\mathcal{G}=(V_{s},\mathcal{E})$, the computational domain $\Omega$ is partitioned into distinct subdomains and each subdomain is allocated to a GPU. Each GPU is allocated a distinct subdomain $\Omega_{r}\subset\Omega$. This arrangement entails dividing $\Omega$ into $n_{\text{procs}}$ (total number of GPUs available for parallelization) subdomains $\Omega_{r}$, with each subdomain featuring an extended overlap of length $l$. In all our runs, unless explicitly mentioned, we set $l=r$ to ensure that nodes at the edges of any given spatial partition have complete kernels This extended overlap is pivotal for ensuring comprehensive kernel construction at the interior edges of the domain, thereby circumventing the issues of incomplete kernel formation that can result in discontinuities in the predicted solutions. This approach is depicted in figure 1
@@ 4.1 Darcy Flow (Structured Data)
We explore an example utilizing a structured grid, specifically focusing on 2-D Darcy flow that describes the fluid flow through a porous medium. 2-D Darcy flow equation on the unit box is a second order elliptic PDE given as:
(˙$-\nabla\dot{(}a(x)\nabla u(x))=f(x)\quad\forall x\in(0,1)^{2}$ (3)
$u(x)=0\quad\forall x\in\partial(0,1)^{2}.$ (4)
Here, $a$ and $f$ are the spatially varying diffusion coefficients and the forcing field, respectively. $u$ is the solution field on the 2-D domain. This example parallels the approach used in the Graph Neural Operator (GNO) as discussed by  (Li et al., 2020a), adapted for a single GPU setup with a notable distinction: we update the edge attributes following each radius hop or message-passing step, and both edges and nodes are subjected to sampling. The experiment investigates the mapping $a\rightarrow u$.
@@
Figure 3: Comparison of model performance between single GPU (S-MPNN) – RMSE=$5.3\times 10^{-6}$, four-GPU (DS-MPNN4) – RMSE=$4.71\times 10^{-6}$, and GCN – RMSE=$4\times 10^{-3}$, for a test sample of Darcy flow dataset.
Table 1 compares the performance of GCN and MPNN models on the Darcy flow dataset, using both single and 4-GPU setups. GCN under-performs relative to MPNN, highlighting the superiority of edge-based methods in PDE surrogate modeling, even for simpler datasets. This discrepancy is largely due to GCN’s omission of edge representation, a key feature in node-based models. Furthermore, the DS-MPNN model shows comparable results on both single and 4-GPU configurations. The minor variance in test error between these two setups might result from the segmentation of the computational graph across different domains. Figure 3 corroborates this observation, illustrating similar predictions from both DS-MPNN and S-MPNN configurations, while GCN lags behind as it excessively smooths the solution and fails to predict the roughness at the boundaries.
@@ 4.2 AirfRANS (Unstructured Data)
The AirfRANS dataset corresponds to the solution of incompressible steady-state Reynolds Averaged Navier Stokes (RANS) equations in 2-D, given by
$\mathbf{U}.\nabla\mathbf{U}=-\frac{1}{\rho}\nabla P+(\nu+\nu_{t})\nabla^{2}\mathbf{U},$ (6)
.$\nabla.\mathbf{U}=0.$ (7)
Here, $\mathbf{U}=[U_{x},U_{y}]$ represents the mean 2-D velocity components, and $P$ denotes the mean pressure. The variable $\nu_{t}$ signifies the kinematic turbulent viscosity, which is spatially varying, while $\nu$ denotes the constant kinematic viscosity. In this dataset, the learning process involves mapping the airfoil shape and angle of attack to $[U_{x},U_{y},P,\nu_{t}]$, as further elaborated in the subsequent sections.
@@ 4.2.1 Low-fidelity AirfRANS dataset
Figure 4(a) illustrates the training loss trajectories of S-MPNN and DS-MPNN4. The convergence patterns, indicated by the overlapping trend in training loss, suggest that DS-MPNN4 achieves a level of training convergence comparable to that of S-MPNN. Furthermore, upon adopting an alternative scheduler strategy—specifically, reducing the learning rate once the training loss plateaus, a similar convergence pattern is observed in figure 4(b). This consistency underscores the training robustness of the DS-MPNN model.
(a) Scheduler: One cycle learning rate
(b) Scheduler: Dropping learning rate on plateau
Figure 4: Training loss vs. epochs for different schedulers.
@@
Table 4: Variation of test RMSE as a function of overlap length ($l$). Kernel radius is set fixed at $r=0.3$.
$l=0$ $l=0.15$ $l=0.3$ $l=0.6$
Test RMSE 0.124 0.096 0.097 0.097
Table 4 presents the test RMSE errors for various extended overlap lengths ($l$) for a fixed kernel radius $r=0.3$. As the overlap radius increases from 00 to $0.15$, the error decreases significantly. However, beyond $0.15$, the variations are minimal. As previously mentioned in section 3, we set $l=r$ in all our experiments unless explicitly mentioned. This is done to ensure that nodes at the edges of any given spatial partition have complete kernels. It should be noted that the kernel radius $r$ is an use-case specific hyperparameter.
@@ 4.2.2 High-fidelity AirfRANS dataset
For the high-fidelity mesh analysis, we employ the standardized AirfRANS dataset as detailed by Bonnet et al. (Bonnet et al., 2022a). Our focus in this study is to show the effectiveness of DS-MPNN across various network hyperparameters rather than to achieve the best accuracy. Hence, we use a specific subset of the original AirfRANS dataset, termed as the ‘scarce’ dataset by the authors, that consists of 180 training samples and 20 test samples. The dataset encompasses a Reynolds number variation between 2 million and 6 million, and the angle of attack ranges from $-5^{\circ}$ to $15^{\circ}$. The input node attributes bear resemblance to the low-fidelity dataset 4.2.1 but with an expanded dimension of unit surface outward-pointing normal for the node $v_{i}^{l={0}}\in\mathbb{R}^{7}$, and the edge attributes are characterized by $e_{ij}^{(l)}\in\mathbb{R}^{10}$. The model output is denoted as $y_{i}\in\mathbb{R}^{4}$. Table 5 presents the test error, quantified as the Root Mean Square Error (RMSE) loss, where the dataset is sampled once. These results indicate that the DS-MPNN framework maintains its efficacy even with a substantial increase in the node count in the dataset, consistently performing well across different GPU configurations. In contrast, the GCN exhibits poorer performance under similar conditions. Additionally, an experiment involving a network configuration without inter-GPU communication, but with distinct domains distributed across GPUs, resulted in divergent training outcomes which are different from (Strönisch et al., 2023) where no communication model worked better. This underscores the critical role of communication between GPUs for maintaining model stability and performance. Table 5 shows that using the DS-MPNN framework with more GPUs increases training and inference speeds. This is because each GPU deals with fewer edges and nodes than it would if the entire domain was on a single GPU, which speeds up model execution. The exploration of scalability and the associated communication overhead for DS-MPNN is comprehensively addressed in Section 5.
@@
5 Ablation Studies on Scalability and Communication Overhead
The subsequent ablation experiments were conducted to evaluate the scalability and communication burden of DS-MPNN4 in comparison with the standard single GPU framework, S-MPNN. These studies were carried out using the ‘scarce’ high-fidelity AirfRANS dataset, which consists of 180 training samples. For these experiments, the kernel radius and the number of nodes were predetermined at $r=0.05$ and $6000$, respectively. Unless specifically mentioned, the baseline overlap length is set to the kernel radius. Figure 6, clearly demonstrate the associated cost of communication for DS-MPNN4. The training and inference costs for communication are higher compared to scenarios with no communication across various non-zero lengths of the overlap region in the 4 sub-domains. Nevertheless, despite the communication overhead, DS-MPNN4 is consistently better than S-MPNN in terms of training and inference duration.
Figure 6: Training time per epoch and inference time as a function of the length of the overlap region ($l$).
We also increase the sampled node count $s$ to ascertain the impact of scalability and communication overhead. This increase in node samples becomes imperative for larger numerical grids. Figure 7 shows the increase in the training speed of DS-MPNN4 compared to S-MPNN. The increase in speed is attributed to the reduction in the time required for graph kernel $\mathcal{G}$ generation within DS-MPNN4, as a consequence of decreased number of edge formations required across each GPU domain. Correspondingly, the percentage of training time spent on communication decreases as the node count increases, because, while the aggregate data for communication grows with the node count, the substantial bandwidth of GPUs maintains a constant communication time.
Figure 7: Performance of DS-MPNN4 against S-MPNN with node increase.
Figure 8: Training and inference time as a function of number of distributed domains.
Moreover, scalability appears in figure 8 that shows a decrease in training and inference time with an increasing number of domains from one (S-MPNN) to four (DS-MPNN4) GPU setup. This is consistent with our previous observations.
@@ 6 Conclusions
We address the issue of scaling edge-based graph methods to larger numbers of nodes by introducing distributed training for message-passing neural networks (MPNNs). When combined with node sampling techniques, this distributed approach allows us to scale a larger number of nodes with losing no or minimal loss in accuracy as compared to the single-GPU implementation while achieving a considerable decrease in training and inference times. We also show comparisons to the graph convolution networks (GCNs) and establish that edge-based methods like MPNNs outperform node-based methods like GCNs - This work opens up new avenues for the use of edge-based graph neural networks in problems of practical interest where the number of nodes can be impractically large, $\sim O(10^{5}-10^{6})$, for single-GPU memory limits.  The deployment of DS-MPNN is complex, notably in managing unstructured grids and their adjacent elements. However, the domain of distributed graph training is advancing swiftly. Notably, libraries tailored for these challenges have been introduced into the latest iteration of PyTorch Geometric (Fey & Lenssen, 2019), version 2.5.0.
@@ 7 Impact statements
Appendix A Training Algorithm
Algorithm 1 Training of DS-MPNN.
Input: Node encoder: $\mathcal{N}_{e}$, Node decoder: $\mathcal{N}_{d}$, Graph kernel: $\mathcal{K}_{\phi}$; Domain: $\Omega$, Sampled-domain: $\Omega_{s}$, Sub-domain: $\Omega_{r}$; grid coordinates and initial conditions: $v^{l={0}}$., Target values: $Y$, No of training samples: $N_{train}$, Number of epochs: $E$, Number of distributed domains: $n_{proc}$.
for $j$ = $1$ to $N_{train}$ do
       $\Omega_{s}\leftarrow rand(\Omega)$triangleright Sample nodes from domain $\Omega$
$\Omega_{r}\leftarrow\Omega_{s}$triangleright Decompose domain based on available gpus
$v^{l={0}}\leftarrow\Omega_{r}$triangleright Grid coordinates and initial values to create nodes
$Y\leftarrow\Omega_{r}$triangleright Get targets in the sub-domain
$edge_{I}\leftarrow edgeindexcreator(\Omega_{r},r)$triangleright Create edge index for kernel radius r
$edge_{i}\leftarrow rand(edge_{I})$triangleright Sample edge index
$e^{l}\leftarrow(dv^{l={0}})$triangleright Create edge attributes ($dv=v_{i}-v_{j})$
$trainloader\leftarrow tuple(v^{l={0}},Y,edge_{i},e^{l})$
for $\textrm{epoch}=1$ to $E$ do
       $\triangleright$ This loop run across $n_{proc}$ GPUs
$v^{l={0}},Y,edge_{i},e^{l}\leftarrow trainloader$
$i_{b}\leftarrow getindex(v^{l={0}})$triangleright Tracking index $i_{b}$ for unstructured grids
for $k$ = $1$ to $h$ do
             if k = 1 then
                   ${v^{l}_{L}}\leftarrow\mathcal{N}_{e}\left({v^{l={0}}},{w}\right)$triangleright Node encoding
            $v_{res}\leftarrow v^{l}_{L}$ $v^{l}_{L}\leftarrow\mathcal{K}_{\phi}\left({v^{l}_{L}},edge_{i},e^{l},{w}\right)$triangleright Graph convolution
$v^{l}_{L}\leftarrow v^{l}_{L}+v_{res}$ $v_{i}^{l}\leftarrow\mathcal{N}_{d}\left(v^{l}_{L},{w}\right)$triangleright Node decoding
$v^{l}_{L}\leftarrow Comm(i_{b},\Omega$,$v^{l}_{L}$) $\triangleright$ Communicate node values in latent space
$v^{l}\leftarrow Comm(i_{b},\Omega,v^{l})$triangleright Communicate output values
$e^{l}\leftarrow dv^{l}$triangleright Update edge attributes with new values
       $\mathcal{L}_{local}$ = $\textrm{MSE}\left({{{v^{l}}_{i_{(int)}}},{Y_{i_{(int)}}}}\right)$triangleright MSE on interior points
$\mathcal{L}=\sum_{i=1}^{n_{proc}}\mathcal{L}_{local}$triangleright Getting loss from all procs
$\nabla{w_{local}}\leftarrow\textrm{Backprop}(\mathcal{L})$triangleright Get local gradients
$\nabla{w}$ = $\sum_{i=1}^{n_{proc}}\nabla{w_{local}}$triangleright Sum local gradients
${w}\leftarrow{w}-\eta\nabla{w}$triangleright Update parameters
Output: Trained DS-MPNN model $\mathcal{N}(\mathcal{N}e,\mathcal{N}d,\mathcal{K}\phi)$